Solve the system of equations using any method:
x/3-3y/4=-1/2
x/6-y/8=3/4
I tried to solve this problem using elimination and graphing methods. It used up 2 whole pieces of paper and I don't think I got the right answer. Which method is best for this problem? Please help.Thank you!

Answer by stanbon(75887)

Solve the system of equations using any method:
x/3-3y/4=-1/2
x/6-y/8=3/4
--------------------
Modify the equations:
Multiply through the 1st Eq. by 12
Multiply through the 2nd Eq. by 24
---------
4x - 9y = -6
4x - 3y = 18
-------------
Subtract the 1st Eq. from the 2nd and solve for "y":
6y = 24
y = 4
---
Substitute into 4x-3y=18 to solve for "x":
4x -3*4 = 18
4x -12 = 18
4x = 30
x = 15/2
----------------------
Cheers,
Stan H.
========================